### v2

# Competitive Equilibrium for Chores: from Dual Eisenberg-Gale to a Fast, Greedy, LP-based Algorithm

Abstract.
We study the computation of competitive equilibrium for Fisher markets with $n$ agents and $m$ divisible chores. Competitive equilibria for chores are known to correspond to the nonzero KKT points of a program that minimizes the product of agent disutilities, which is a non-convex program whose zero points foil iterative optimization methods. We introduce a dual-like analogue of this program, and show that a simple modification to our “dual” program avoids such zero points, while retaining the correspondence between KKT points and competitive equilibria. This allows, for the first time ever, application of iterative optimization methods over a convex region for computing competitive equilibria for chores. We next introduce a greedy Frank-Wolfe algorithm for optimization over our program and show a new state-of-the-art convergence rate to competitive equilibrium. Moreover, our method is significantly simpler than prior methods: each iteration of our method only requires solving a simple linear program. We show through numerical experiments that our method is extremely practical: it easily solves every instance we tried, including instances with hundreds of agents and up to 1000 chores, usually in 10-30 iterations, is simple to implement, and has no numerical issues.
1 Introduction
We study competitive equilibrium (CE) for Fisher markets with chores, where a set of divisible chores need to be allocated among a set of agents. Each agent incurs disutility from the chores assigned to her, and is compensated with payments for the chores she undertakes. We study the most classical utility model, where the disutility of each agent is linear in the chores assigned to her. Each agent is assumed to have a fixed earning requirement. CE is a central solution concept in microeconomics which specifies how markets can assign prices and allocations in a way that satisfies demand and supply. In the case of Fisher markets with chores, prices and allocations form a CE if (i) each agent gets her disutility-minimizing bundle that achieves her earning requirement, and (ii) all chores are fully allocated.
A classical motivation for the Fisher market model with goods, is the competitive equilibrium with equal incomes (CEEI), which is a mechanism for fairly and efficiently dividing a set of goods among a set of agents (Varian 1974). The CEEI idea naturally extends to the case of chores, and still provides similar fairness and efficiency properties. Examples of fair chore allocation problems include family members dividing household chores, workers dividing job shifts, faculty dividing teaching loads, peer review assignment, and so on.
In a seminal paper, Bogomolnaia et al. (2017) showed that CE exists in chores Fisher markets with 1-homogeneous convex disutility functions [^1] Bogomolnaia et al. (2017) measure utility, in which case the utility function is concave; for our purposes it is easier to work with the equivalent disutility of an agent, in which case the corresponding function is convex., and gave an optimization characterization of the set of CE. They also showed that, despite the resemblance to CE with goods, CE with chores has surprisingly different structural and computational properties. Firstly, the set of CE can be disconnected (Bogomolnaia et al. 2017), in contrast to CE with goods where all equilibria form a convex set. Secondly, the CE set for goods is the set of optimal solutions to a convex program known as the Eisenberg-Gale convex program, which is a program that maximizes the product of agent utilities. In the case of chores, the set of CE is still captured by an analogous program which attempts to minimize the product of agent disutilities. Bogomolnaia et al. (2017) show that every KKT point of this program (call it the EG program for chores) where every agent receives strictly positive distuility corresponds to a CE. Importantly, KKT points where some agent receives zero disutility do not correspond to CE, yet such points are the global minimizes of the program. This effectively means that EG for chores has an open constraint, and the global infimum is achieved at the closure of this open constraint. This is highly problematic from an optimization perspective, because these zero points may attract iterative optimization methods. Due to this algorithmic difficulty, polynomial-time algorithms are only known for computing approximate CEEI (Boodaghians et al. 2022, Chaudhury et al. 2022), and existing methods all rely on highly specialized subroutines that are not very practical.
As is the case for the EG program for goods, we will work with the logarithm of the product of agent disutilities, which converts the product of disutilities into a concave function, albeit one that we need to minimize. Note that the domain of the logarithm function automatically rules out allocations with zero disutility for some agent, and thus all KKT points of this objective correspond to CE. However, this does not solve the problem of applying iterative optimization methods: under this formulation the objective has infeasible points on the boundary, corresponding to zero disutility allocations, such that the objective converges to negative infinity as we approach these points. We refer to these points as poles, in analogy to their usage in complex analysis. In this formulation, starting from almost every feasible points, standard iterative optimization methods move towards these poles and the objective tends to negative infinity (Boodaghians et al. 2022).
To circumvent the poles issue, sophisticated and highly specialized iterative methods have been developed (Boodaghians et al. 2022, Chaudhury et al. 2022) for finding a nonzero KKT point. However, both methods require exactly solving a non-linear convex program at each iteration – the projection subroutine in (Boodaghians et al. 2022), and the balanced flow subroutine in (Chaudhury et al. 2022). Exactly solving a nonlinear convex program at every iteration can be expensive, and thus these methods are not practical for large-scale problems. A major roadblock for developing more practical algorithms is the presence of poles in the EG for chores program. This opens up a natural question:
Does there exists an optimization formulation of the chores problem that avoids the poles issue?
An affirmative answer to this question would open up the CE for chores problem to more standard iterative optimization methods, making it more tractable both theoretically and practically. As we discuss below, we show that such a formulation indeed exists, and it leads to significantly faster algorithms.
Organization.
The remainder of this section discusses our contributions and related work. Section 2 gives preliminaries for Fisher markets with chores and existing mathematical programs. Section 3 introduces two new (dual EG) programs capturing CE with chores. Section 4 provides with a new greedy, fast, LP-based algorithm to compute a CE, and proves its convergence properties. Section 5 demonstrates practical efficiency of our new algorithm and Section 6 includes conclusion and discussion.
1.1 Our Contributions
Novel program with no poles.
As our first main result, we give a novel program that maximizes a convex function subject to simple linear constraints. We show that our program circumvents the issue of poles of the EG program for chores, while maintaining the one-to-one correspondence between CE and KKT points. To the best of our knowledge, this is the first program that makes the chores problem amenable to standard gradient-descent-type algorithmic approaches.
The inspiration for our new program comes from the dual of the EG program in the context of goods. The dual EG program for goods has been used extensively for deriving effective algorithms, as we discuss in Section 1.2. In a formal sense, there is no duality theory for the chores setting, since we are dealing with nonconvex programs. Nonetheless, we develop a natural “dual” formulation of the EG program for chores. However, just as the primal EG program for chores, the natural dual EG program for chores also has poles. Surprisingly, we show that the poles of our new program can be removed by adding a linear constraint. This constraint not only gets rid of all poles, but also preserves all KKT points without introducing any new KKT points. We call the new program (Chores Dual Redundant). We observe that, although a non-convex optimization problem, (Chores Dual Redundant) has strong duality-like structural properties: its objective matches that of the primal EG program at the KKT points.
Greedy Frank Wolfe algorithm.
Armed with this new program, it is natural to attempt a first-order method to compute a CE with chores. In this paper, we adopt the greedy variant of the famous Frank-Wolfe (FW) method (Frank and Wolfe 1956). In Section 4, we show that this method converges to an $\epsilon$-approximate CE (formally defined in Definition (2)) in $\mathcal{\tilde{O}}(\frac{n}{\epsilon^{2}})$ iterations, matching the convergence guarantees in Boodaghians et al. (2022), Chaudhury et al. (2022), in terms of dependence on $\epsilon$, while improving the dependence on the number of buyers and chores. To be specific, the exterior point method by Boodaghians et al. (2022) requires $\mathcal{\tilde{O}}(\frac{n^{3}}{\epsilon^{2}})$ iterations, and the combinatorial algorithm by Chaudhury et al. (2022) requires $\mathcal{\tilde{O}}(\frac{nm}{\epsilon^{2}})$ iterations. We emphasize that, unlike the algorithms in Boodaghians et al. (2022), Chaudhury et al. (2022), an iteration of GFW involves solving only a single LP, as opposed to more general convex programs in prior methods. Further, the dependence on problem size in the bound on the number of iterations for our GFW method is significantly smaller than prior methods (Boodaghians et al. 2022, Chaudhury et al. 2022). Combining these facts, we expect GFW to be considerably faster in practice, especially for large instances. We confirm this intuition through a comprehensive set of empirical results.
Convergence for relative strongly convex maximization.
Consider a problem of maximizing a relative-strongly-convex function $f$ (w.r.t. some reference function $h$) over a nonempty closed convex set. We show that after $T\in\mathcal{O}\Big{(}\frac{f^{*}-f^{0}}{\alpha\epsilon}\Big{)}$ [^2] Here, $f^{*}$ and $f^{0}$ are the optimal and initial objective values, respectively. $\alpha$ denotes the modulus of the relatively strong convexity. iterations of the GFW algorithm, there exists an iterate $k\in[T]$, such that the distance between $x_{k}$ and $x_{k-1}$ is small compared to $\epsilon$, as measured by the Bregman divergence induced by the reference function. See Section 4.1 for our generalized results. We use this result, along with a particular choice of $h$, to show that GFW finds an $\epsilon$-approximate CE. We believe that this general approach could lead to interesting guarantees for other problems of maximizing a convex function (e.g. the approximate $\ell_{0}$-penalized PCA [^3] Principal component analysis. problem (Luss and Teboulle 2013), multiplicative problems (Benson 1995), etc.). In the appendix, we also present an alternative primal-dual analysis of GFW specifically for the (Chores Dual Redundant) program, with mildly stronger guarantees (see Appendix A).
Empirical evaluation.
We compare our GFW method to the state-of-the-art existing method: the exterior-point method (EPM) from Boodaghians et al. (2022), on a large number of instances sampled from five different classes of random disutility distributions, as well as on a semi-synthetic dataset derived from a conference review bidding dataset. Our GFW algorithm solves every instance that we tried it on, usually in 10-30 iterations, and in less than a minute for instances with several hundred buyers and chores. In contrast, the prior EPM fails on a large number of instances (for some disutility distributions it fails on every instance once the number of buyers and chores is on the order of a few hundred each), though it is usually competitive with our method in the cases where it does not fail. We conclude that our GFW method is the first highly practical method: it is extremely robust and simple to implement, and can computes an exact CE within a minute for every instance we tried, even with upwards of 300 buyers and chores.
1.2 Related Work
CE with goods has seen a long line of research since the 1950s. Early work (Arrow and Debreu 1954) showed the existence of CE, and also formulated convex programs (Eisenberg and Gale 1959) that capture all CE under specific utility functions. In particular, Eisenberg and Gale (1959) show that any allocation that maximizes the product of the utility functions of the agents corresponds to a CE. Despite the existence of the convex program, and polynomial bounded rationality of the optimal solutions, there has been a long study on a variety of algorithms to compute CE under linear utilities, including interior point methods (Jain 2007, Ye 2008), combinatorial methods (Devanur et al. 2008, Orlin 2010, Duan and Mehlhorn 2015, Garg and Végh 2019, Chaudhury and Mehlhorn 2018), and first-order methods and dynamics (Birnbaum et al. 2011, Zhang 2011, Gao and Kroer 2020). One of the reasons for designing special-purpose algorithms is attributed to the lack of practically fast convex program solvers for large instances (Vazirani 2012)– for instance, convex program solvers are significantly slower than LP solvers, and are practical only for medium-size instances (see benchmarks in Mittelmann). Furthermore, the broad algorithmic literature has also revealed several structural and economically valuable insights on CE (Duan et al. 2016, Devanur et al. 2013, Garg and Végh 2019). Finally, CE has been used in a variety of applications, especially in the case of linear utility or disutility functions (Goldman and Procaccia 2015, Conitzer et al. 2022, Allouah et al. 2023). For example, the online platform Spliddit has seen tens of thousands of users for the past few years (Goldman and Procaccia 2015). Spliddit uses additive preferences which is the parallel of linear utility functions in the context of sharing indivisible items. CE with linear utilities has also seen extensive use in online fair allocation (Azar et al. 2016, Banerjee et al. 2022, Gao et al. 2021, Benadè et al. 2023).
The problem of computing a CE involving chores has attracted significant attention since the characterization result of Bogomolnaia et al. (2017). There are polynomial-time algorithms if the number of agents or the number of chores is a constant: in this special case, Brânzei and Sandomirskiy (2023) showed CE with chores can be computed in strongly polynomial time and Garg and McGlaughlin (2020) showed a polynomial-time algorithm for computing CE in settings that include both goods and chores. Chaudhury et al. (2021) developed a simplex-like algorithm based on Lemke’s scheme for additively separable piecewise linear concave utilities, and showed that it can be implemented for small instances with 10-20 agents and chores. As mentioned previously Boodaghians et al. (2022) and Chaudhury et al. (2022) developed polynomial-time algorithms for computing approximate CE with chores, where their iteratations require solving non-linear convex programs.
2 Preliminaries
2.1 Fisher Market with Chores
Analogous to Fisher markets in the goods setting, a chores market comprises of a set of $n$ agents, and a set of $m$ divisible chores. Each agent $i$ has a disutility of $d_{ij}$ for one unit of chore $j$. Given a bundle of chores $x_{i}\in\mathbb{R}^{m}_{\geq 0}$, where $x_{ij}$ denotes the amount of chore $j$, agent $i$’s total linear disutility for bundle $x_{i}$ is $\sum_{j\in[m]}d_{ij}x_{ij}$. We also write $\sum_{j\in[m]}d_{ij}x_{ij}$ as $\langle d_{i},x_{i}\rangle$, where $d_{i}=(d_{i1},d_{i2},\dots,d_{im})$. Further, each agent $i$ has an earning requirement of $B_{i}\geq 0$ units of money.
In a CE for a chores market, we wish to find a price $p_{j}$ for each chore $j$, and an allocation $x\in\mathbb{R}^{n\times m}_{\geq 0}$ of chores to the agents, where each $x_{ij}$ is the amount of chore $j$ allocated to agent $i$, such that (i) each agent $i$ gets a disutility minimizing bundle $x_{i}=(x_{i1},x_{i2},\dots,x_{im})$, subject to her earning requirement of $B_{i}$ units of money, and (ii) all chores are allocated. Formally,
Definition 1.
A price vector $p\in\mathbb{R}^{m}_{\geq 0}$, and an allocation $x\in\mathbb{R}^{n\times m}_{\geq 0}$ satisfy competitive equilibrium (CE) if and only if
  • (E1)
    $\langle p,x_{i}\rangle=B_{i}$ for all $i\in[n]$ ;
  • (E2)
    $\langle d_{i},x_{i}\rangle\leq\langle d_{i},y_{i}\rangle$ for all $y_{i}$ such that $\langle p,y_{i}\rangle\geq\langle p,x_{i}\rangle$, for all $i\in[n]$ ;
  • (E3)
    $\sum_{i}x_{ij}=1$ for all $j\in[m]$
Condition (E1) and (E2) ensure that each agent gets a disutility-minimizing bundle subject to her earning constraint, and (E3) captures that all chores are allocated in this process. CEEI is one of the most well-studied special cases of CE, where $B_{i}=1$, for all $i\in[n]$. It is a well known (see Bogomolnaia et al. (2017, Lemma 1)) that if $(p,x)$ satisfy CEEI, then (i) $x$ is envy-free, i.e., $d_{i}(x_{i})\leq d_{i}(x_{j})$ for all $i,j\in[n]$, and (ii) $x$ is Pareto-optimal, i.e., there is no $y$ such that $d_{i}(y_{i})\leq d_{i}(x_{i})$ with at least one strict inequality. CEEI satisfies many other desirable properties like core-stability, and we refer the reader to Bogomolnaia et al. (2017) for a more detailed explanation of the remarkable fairness and efficiency properties satisfied by CEEI. The existence of CE in the chores market can be shown by adapting the fixed point arguments in Arrow and Debreu (1954).
Our algorithms will be concerned with computing approximate CE solutions. To that end, we relax conditions (E1), (E2), and (E3) in Definition 1 as follows,
Definition 2 ($\epsilon$-approximate CE (Boodaghians et al. 2022)).
A price vector $p\in\mathbb{R}^{m}_{\geq 0}$, and an allocation $x\in\mathbb{R}^{n\times m}_{\geq 0}$ satisfy $\epsilon$-competitive equilibrium (CE) if and only if
  • (1)
    $(1-\epsilon)B_{i}\leq\langle p,x_{i}\rangle\leq\frac{1}{1-\epsilon}B_{i}$ for all $i\in[n]$;
  • (2)
    $(1-\epsilon)\langle d_{i},x_{i}\rangle\leq\langle d_{i},y_{i}\rangle$ for all $y_{i}$ such that $\langle p,y_{i}\rangle\geq\langle p,x_{i}\rangle$, for all $i\in[n]$;
  • (3)
    $1-\epsilon\leq\sum_{i}x_{ij}\leq\frac{1}{1-\epsilon}$ for all $j\in[m]$.
$\epsilon$-approximate CEEI is also analogously defined. Furthermore, we consider a stronger notion of approximate CE, called $\epsilon$-strongly approximate CE.
Definition 3.
A pair of $(p,x)$ is said to be an $\epsilon$-strongly approximate CE if $(E2)$ and $(E3)$ holds strictly and only $(E1)$ is relaxed to $(1)$.
2.2 Mathematical Programming Approaches to Competitive Equilibria
In the case of CE for goods, it is known that a CE allocation can be obtained from the Eisenberg-Gale convex program. In the goods case with linear utilities, each buyer has a value vector $v_{i}\in\mathbb{R}_{+}^{m}$ describing their value per good, and they compute demands by maximizing their obtained value subject to budget constraints. The EG program for this case is stated on the left below.
[…]
On the right in (1) is the dual of the EG program, which minimizes a function of $p$ and $\beta$, where $p$ and $\beta$ can be interpreted as prices for the goods, and inverse bang-per-buck variables for the buyers, respectively. We will use the dual of the goods EG program as inspiration for our chores approach.
Bogomolnaia et al. (2017) show an interesting characterization of all CE in the chores market. Similar to the case of goods, they show that for chores there is also an EG-style program for computing a CE. More specifically, they show that there is a one-to-one correspondence between the set of CE and all KKT points of the following program,
$\inf_{x\geq 0}$ $\prod_{i\in[n]}\bigg{(}\sum_{j\in[m]}d_{ij}x_{ij}\bigg{)}^{B_{i}}$ (2)
s.t. $\sum_{i\in[n]}x_{ij}=1\quad\;\;\;\;\;\forall\,j\in[m]$
$\sum_{j\in[m]}d_{ij}x_{ij}>0\quad\forall\,i\in[n].$
Applying the logarithm, which is a monotone transformation, to the objective yields a mathematical program analogous to the EG program for goods,
$\inf_{x\geq 0}$ $\sum_{i\in[n]}B_{i}\cdot\log\bigg{(}\sum_{j\in[m]}d_{ij}x_{ij}\bigg{)}$ (3)
s.t. $\sum_{i\in[n]}x_{ij}=1\quad\;\;\;\;\;\forall\,j\in[m].$
After applying the logarithm, we no longer need an explicit open constraint, since the domain constraint on the logarithm already enforces strictly positive disutilities.
However, this program has several disadvantages compared to the goods case. Most notably, it is a non-convex program, and the objective function can tend to negative infinity within the feasible region (as the disutility of an agent tends to zero). This makes it challenging algorithmically, because iterative methods that operate inside the feasible set are attracted to these “negatively-infinite-objective” points (Boodaghians et al. 2022).
2.3 Bregman Divergence and Generalized Strong Convexity
Let $h:\Omega\rightarrow\mathbb{R}$ be a differentiable, strictly convex function defined on a convex set $\Omega$. The Bregman divergence associated with $h$ between $y,x\in\Omega$ is defined as $D_{h}(y\,\|\,x)=h(y)-h(x)-\langle\nabla h(x),y-x\rangle$. The function $h$ is called the reference function of $D_{h}(\cdot\,\|\,\cdot)$. For example, if $h(x)=\frac{1}{2}{\left\|x\right\|}^{2}$ then $D_{h}(y\,\|\,x)=\frac{1}{2}{\left\lVert y-x\right\|}^{2}$. Since $h$ is strictly convex, $D_{h}(y\,\|\,x)\geq 0$ for all $y,x\in\Omega$, and $D_{h}(y\,\|\,x)=0$ iff $y=x$.
We say a differentiable function $f$ is $\alpha$-strongly convex with respect to $h$ if $\text{dom}\,f\subseteq\Omega$ and
$f(y)\geq f(x)+\langle\nabla f(x),y-x\rangle+\alpha D_{h}(y\,\|\,x)\quad\forall\,y,x\in\text{dom}\,f.$ (4)
We refer Eq. 4 with an arbitrary $D_{h}(\cdot\,\|\,\cdot)$ as generalized strong convexity or relatively strong convexity, and it reduces to (standard) strong convexity if $D_{h}(y\,\|\,x)=\frac{1}{2}{\left\lVert y-x\right\|}^{2}$.
3 The Chores Dual Program
In this Section, we introduce a novel program with no poles, such that there is a one-to-one correspondence between CE in the chores market and the KKT points of our program. Our inspiration for this program comes from the dual program for CE in the goods setting (Eq. 1). In the goods case, the dual program has been used extensively in order derive various algorithmic approaches for computing CE. The proportional response dynamics can be derived by a change of variable and dualization step on the goods EG dual (Zhang 2011, Birnbaum et al. 2011), and the PACE algorithm for online Fisher market equilibrium is also derived from this program (Gao et al. 2021, Liao et al. 2022). Given the algorithmic usefulness of the EG dual in the goods case, it is natural to ask whether there is an analogue for the chores case. The EG program for chores, however, is not a convex program, and thus convex duality does not let us directly derive a dual. Nonetheless, we may guess that the chores EG “dual” should be the maximization version of the goods EG dual, analogously to how the chores “primal” is the minimization version of the goods EG primal. This guess yields the following program which maximizes a convex function:
$\sup_{\beta\geq 0,p\geq 0}$ $\sum_{j}p_{j}-\sum_{i}B_{i}\log(\beta_{i})$ (Chores Dual)
s.t. $\;p_{j}\leq\beta_{i}d_{ij}\quad\quad\forall\;i,j.$
Henceforth we shall refer to this as the chores dual, even though it is not a dual program in the formal sense of convex duality. It is worth noting that, for example, we do not have global weak duality: some feasible pairs $(\beta,p)$ of (Chores Dual) yield a dual objective that does not lower bound the primal objective (this is easy to see, since the dual objective tends to infinity near poles).
Even though it is not a formal dual of the primal program, it turns out to behave in a dual-like manner at KKT points. Indeed, the KKT points of this program yield chores CE, similar to how the primal EG for chores yields CE at all KKT points. We emphasize here that in much of the prior literature on Nash welfare minimization for chores, the focus has been on finding non-zero-disutility KKT points; once we take the log of the Nash welfare objective, this constraint is implicitly enforced as a domain constraint on the log function, thus all KKT points become CE, both in the “primal” and “dual” EG programs for chores.
Theorem 1.
There is a one-to-one correspondence between competitive equilibria of a chores Fisher market and the KKT points of (Chores Dual).
Proof.
First, we show that any KKT point of (Chores Dual) is a CE. Let $(\tilde{p},\tilde{\beta})$ denote a feasible point in the primal space of (Chores Dual), and $\tilde{x}$ be the dual variable associated with the constraint in (Chores Dual). Note that we overload notation slightly here: $(p,\beta,x)$ can, at this stage, only be understood to be primal and dual variables that arise from the programs, whereas we previously used these notations to refer to market quantities like prices and allocations. We will show that the primal and dual variables do in fact have these interpretations, thus justifying the notation. By the KKT conditions, any KKT point $(\tilde{p},\tilde{\beta},\tilde{x})$ of (Chores Dual) satisfies
  1. (i)
    $p,x\geq 0,\beta>0$;
  2. (ii)
    $d_{ij}\beta_{i}\geq p_{j}$ and $x_{ij}\left(d_{ij}\beta_{i}-p_{j}\right)=0$ for all $i\in[n],j\in[m]$;
  3. (iii)
    $\sum_{i}x_{ij}\geq 1$ and $p_{j}\left(\sum_{i}x_{ij}-1\right)=0$ for all $j\in[m]$;
  4. (iv)
    $\sum_{j}d_{ij}x_{ij}\geq\frac{B_{i}}{\beta_{i}}$ and $\beta_{i}\left(\sum_{j}d_{ij}x_{ij}-\frac{B_{i}}{\beta_{i}}\right)=0$ for all $i\in[n]$.
Note the $\beta>0$ (instead of $\beta\geq 0$) in $(i)$ above. This is implied by the domain constraint of the log function on $\beta_{i}$s. By (ii), we have $\frac{\tilde{p}_{j}}{d_{ij}}=\tilde{\beta}_{i}\geq\frac{\tilde{p}_{j^{\prime}}}{d_{ij^{\prime}}}$ for all $j^{\prime}\in[m]$, if $\tilde{x}_{ij}>0$ for all $i\in[n],j\in[m]$, which is equivalent to $(E2)$ in Definition 1. Combining (ii) and (iv), we have $\sum_{j}\tilde{p}_{j}\tilde{x}_{ij}=\sum_{j}d_{ij}\tilde{\beta}_{i}\tilde{x}_{ij}=B_{i}$ for all $i\in[n]$. Also, (iii) leads to $\sum_{i}\tilde{x}_{ij}=1$ if $\tilde{p}_{j}>0$ for all $j\in[m]$. Thus, by definition, $(\tilde{p},\tilde{x})$ constitutes a CE if we regard $p$ as price and $x$ as allocation.
Next, we prove that any CE is a KKT point of (Chores Dual). Given any CE $(p,x)$, we denote $\beta_{i}:=\frac{B_{i}}{\sum_{j}d_{ij}x_{ij}}$ for all $i\in[n]$. Then, $(p,\beta,x)$ trivially satisfies (i), (iii) and (iv). Since, for all $i$, $\sum_{j}d_{ij}x_{ij}=\max_{x^{\prime}_{i}:\langle p,x^{\prime}_{i}\rangle\geq B_{i}}\sum_{j}d_{ij}x^{\prime}_{ij}$ and $\sum_{j}d_{i}jx_{i}j\leq\frac{d_{ij^{\prime}}}{p_{j}}\cdot B_{i}$ for all $j^{\prime}$, we have $d_{ij}\beta_{i}=\frac{d_{ij}B_{i}}{\sum_{j}d_{ij}x_{ij}}\geq\frac{d_{ij}B_{i}}{d_{ij}({B_{i}}/{p_{j}})}=p_{j}$ for all $i\in[n],j\in[m]$. For all $i\in[n]$, $\sum_{j\in[m]}x_{ij}\left(p_{j}-d_{ij}\beta_{i}\right)=B_{i}-B_{i}=0$ since $\langle p,x_{i}\rangle=B_{i}$. Then, (ii) follows since $x_{ij}\left(p_{j}-d_{ij}\beta_{i}\right)\leq 0$ for all $i\in[n],j\in[m]$. Hence, $(p,\beta,x)$ satisfies KKT conditions. ∎
Next we look at another property typically satisfied by primal and dual convex programs: strong duality implies that optimal primal and dual solutions have the same objective. As we already noted, our (Chores Dual) does not even satisfy weak duality. Yet, it turns out that we do have a local form of strong duality, in the sense that the objectives are equal at KKT points. To see this, we next write more explicit versions of the primal and dual, in a way that keeps various important constants around for showing this equality (these same constants are needed in the goods setting).
Consider the primal and dual problems of chores EG in the following forms:
$\inf_{x\geq 0,u\geq 0}$ $f^{p}(x,u):=\sum_{i}B_{i}\log(u_{i})$ (5)
s.t. $\langle d_{i},x_{i}\rangle\leq u_{i}\quad\hskip 20.0pt\forall\;i\in[n]$ $\quad(\beta_{i})$
$-\sum_{i}x_{ij}\leq-1\quad\;\forall\;j\in[m]$ $\hskip 20.0pt(p_{j})$
and
$\sup_{\beta\geq 0,p\geq 0}$ $f^{d}(\beta,p):=\sum_{j}p_{j}-\sum_{i}B_{i}\log{\beta_{i}}+\sum_{i}\big{(}B_{i}\log{B_{i}}-B_{i}\big{)}$ (6)
s.t. $p_{j}\leq d_{ij}\beta_{i}\quad\quad\quad\forall\;i\in[n],\,j\in[m]\hskip 30.0pt(x_{ij}).$
Theorem 2.
If $(x,u,\beta,p)$ is a KKT point [^4] Any pair of primal and dual variables which satisfies the KKT conditions is called a KKT point. of (5), then $(\beta,p,x)$ is a KKT point of (6). Reversely, if $(\beta,p,x)$ is a KKT point of (6) and $u_{i}=\langle d_{i},x_{i}\rangle$ for all $i\in[n]$, then $(x,u,\beta,p)$ is a KKT point of (5). Furthermore, if $(x,u,\beta,p)$ is a KKT point of (5), then
$f^{p}(x,u)=f^{d}(\beta,p).$
Proof.
The first part is obvious due to the one-to-one correspondence between KKT points and CE for both programs. Since $(x,u,\beta,p)$ is a KKT point of (5), we have $\beta_{i}=\frac{B_{i}}{u_{i}}$ for all $i\in[n]$. Since $(p,\beta,x)$ is a KKT point of (6), and thus a CE of the corresponding chores Fisher market, we have $\sum_{j}p_{j}=\sum_{j}\sum_{i}p_{j}x_{ij}=\sum_{i}\sum_{j}p_{j}x_{ij}=\sum_{i}B_{i}$. By these, we can obtain that
$\sum_{j}p_{j}-\sum_{i}B_{i}\log{\beta_{i}}+\sum_{i}\big{(}B_{i}\log{B_{i}}-B_{i}\big{)}=\sum_{i}B_{i}\log{\frac{B_{i}}{\beta_{i}}}+\sum_{j}p_{j}-\sum_{i}B_{i}=\sum_{i}B_{i}\log{u_{i}}.$
Similarly to the primal EG for chores, the dual EG for chores has problems from an optimization perspective. One might try to find a KKT point by finding a local maximum of (Chores Dual) using some sort of iterative optimization method on the interior of the convex feasible region. However, this runs into trouble because (Chores Dual) has two ways that the objective tends to infinity, generally because we can either let $\beta_{i}$ tend to zero in order for $-\log\beta_{i}$ to tend to infinity (we refer to these as poles, as in the primal case), or by letting the prices tend to infinity, since their contribution to the objective grows faster than $-\log\beta_{i}$.
In Fig. 1(a), we show a diagram which illustrates how naive gradient descent methods fail because they are either attracted to poles, or move along an unbounded direction. This instance is a Fisher market with $2$ agents and $1$ chore. In this instance, $B_{1}=B_{2}=1$, $d_{11}=2$ and $d_{21}=1$. Obviously, there is only one KKT point / CE in this instance: $p_{1}=2$; $\beta_{1}=1,\beta_{2}=2$. As we can see, even starting from some feasible points that are very close to the KKT point, trajectories generated by the gradient ascending algorithm are attracted by either of those poles.
Figure 1: An instance of a Fisher market with $2$ agents and $1$ chore. The gray region (polyhedral) is the feasible region of (Chores Dual). The (only) KKT point / CE ($(\beta_{1},\beta_{2},p_{1})=(1,2,2)$) is marked by the yellow point. The blue arrow denotes the direction of gradient at the KKT point. In the left, two green trajectories are iterates generated by the gradient ascending algorithm, starting from two feasible points. In the right, the blue region (2-dimensional polyhedral) is the feasible region of (Chores Dual Redundant).
Next, we show what we found to be a highly surprising fact: the (Chores Dual) can be “fixed” such that both poles and unbounded directions are avoided, without removing any KKT points. This is achieved by adding a new constraint, which is redundant from a CE perspective. It is easy to see that in any CE, we must have that every buyer earns their budget exactly (otherwise they could do fewer chores and still satisfy their earning constraint), and since every item is allocated exactly, we must have that $\sum_{j}p_{j}=\sum_{i}B_{i}$. Adding this constraint to (Chores Dual) yields the following program:
$\max_{\beta\geq 0,p\geq 0}$ $\sum_{j}p_{j}-\sum_{i}B_{i}\log(\beta_{i})$ (Chores Dual Redundant)
s.t. $\quad\;\;p_{j}\leq\beta_{i}d_{ij}\quad\forall i,j$
$\sum_{j}p_{j}=\sum_{i}B_{i}$
The new constraint $\sum_{j}p_{j}=\sum_{i}B_{i}$ in (Chores Dual Redundant) cuts off directions with infinite objective: 1) prices are upper bounded by $\sum_{i}B_{i}$, so the rays where prices go to infinity are avoided, 2) each $\beta_{i}$ is lower bounded by $\max_{j}p_{j}/d_{ij}\geq\sum_{i}B_{i}/(m\cdot\max_{j}d_{ij})$ since $p_{j}\geq\sum_{i}B_{i}/m$ for some chore $j$. The following lemma formalizes this fact; the proof follows from the above discussion.
Lemma 1.
The optimal value of (Chores Dual Redundant) is at most
$\Big{(}\sum_{i}B_{i}\Big{)}\Big{(}1-\log(\sum_{i}B_{i})+\log(m\max_{i,j}d_{ij})\Big{)}.$
Crucially, it turns out that the new constraint $\sum_{j}p_{j}=\sum_{i}B_{i}$ is redundant, in the sense that its Lagrange multiplier is zero at every KKT point. The redundancy also means that this constraint is implied by the KKT conditions of (Chores Dual), which implies that the one-to-one correspondence between CE and KKT points still holds for (Chores Dual Redundant).
Theorem 3.
There is a one-to-one correspondence between CE of the chores Fisher market and the KKT points of (Chores Dual Redundant).
Proof.
Here, we refer to the two constraints in (Chores Dual Redundant) as the optimal disutility constraints (the first constraints) and the redundant constraint (the second constraint). Let $x$ and $\mu$ be the dual variables corresponding to the optimal disutility constraint and redundant constraint, respectively. By the KKT conditions, $(p,\beta,x,\mu)$ is a KKT point of (Chores Dual Redundant) if and only if it satisfies the conditions (i), (ii) and (iv) in Theorem 1 as well as
  1. (i)
    $\sum_{j}p_{j}=\sum_{i}B_{i}$;
  2. (ii)
    $\sum_{i}x_{ij}+\mu\geq 1$ and $p_{j}\left(\sum_{i}x_{ij}+\mu-1\right)=0$ for all $j\in[m]$.
First, we show that any KKT point of (Chores Dual Redundant) is a CE. To do this, we only need to show $\mu=0$ if $(p,\beta,x,\mu)$ is a KKT point of (Chores Dual Redundant), since the conditions (ii) and (iv) guarantee that every agent chooses her optimal bundle.
Note that
$\sum_{j}p_{j}(1-\sum_{i}x_{ij})=\sum_{i}B_{i}-\sum_{j}\sum_{i}p_{j}x_{ij}=0,$ (7)
since $\sum_{j}p_{j}=\sum_{i}B_{i}$ by the redundant constraint and $\sum_{j}p_{j}x_{ij}=B_{i}$ for all $i$ by the optimal bundle condition. By KKT condition (ii), we have
$0=\sum_{j\in[m]}p_{j}(1-\sum_{i}x_{ij}-\mu)=\sum_{j\in[m]}p_{j}(1-\sum_{i}x_{ij})-\mu\sum_{j\in[m]}p_{j}=-\mu\sum_{i\in[n]}B_{i}.$ (8)
This equality holds if and only if $\mu=0$.
Next, we prove that any CE is a KKT point of (Chores Dual Redundant). Set $\mu=0$ and the rest follows similarly to the second part of the proof of Theorem 1. ∎
In Fig. 1(b) we show how the redundant constraint restricts the feasible set in the instance in Fig. 1(a) in a way that avoids poles.
Because (Chores Dual Redundant) does not have poles (Lemma 1), it allows the use of first-order methods and interior-point methods on the convex polyhedron
$Y=\Big{\{}(\beta,p)\in\mathbb{R}^{n+m}_{\geq 0}:\,p_{j}\leq\beta_{i}d_{ij},\;\forall\,i\in[n],j\in[m],\;\sum_{j\in[m]}p_{j}=\sum_{i\in[n]}B_{i}\Big{\}}.$
This is surprising because past work (Bogomolnaia et al. 2017, Boodaghians et al. 2022) has repeatedly stressed that a challenge for computing CE for chores is that iterative methods that work in the interior are attracted to non-CE points. Our new program completely circumvents this issue. The remainder of the paper will be concerned with developing a provably fast and highly practical iterative method for computing (approximate) KKT points of (Chores Dual Redundant). We note, though, that (Chores Dual Redundant) opens the door to many other optimization methods, and we hope, for example, that efficient interior-point methods can be developed for it, as well as new first-order methods for large-scale problems.
4 Greedy Frank-Wolfe Method
In this section we introduce a simple first-order optimization method for computing chores CE based on (Chores Dual Redundant). The method only requires solving an LP at every iteration, and we will see that it has many desirable properties.
As above, let $Y$ be the convex polyhedron of the feasible set of (Chores Dual Redundant), and let $y=(\beta,p)\in Y$ denote elements of $Y$. Let $f$ be the (Chores Dual Redundant) objective (after dropping $\sum_{j}p_{j}$, which sums to the constant $\sum_{i}B_{i}$), i.e.,
$f(y)=-\sum_{i}B_{i}\log\beta_{i}.$
The algorithm we will use is a greedy variant of the Frank-Wolfe (FW) algorithm (Frank and Wolfe 1956). For a general differentiable minimization problem over a closed convex set, say $\min_{x\in\mathcal{X}}\tilde{f}(x)$, the FW method conducts the following steps iteratively. In every iteration $t$, the FW algorithm finds a minimizer $s_{t}$ to the linear approximation of the function $\tilde{f}$ around the previous iterate point $x_{t-1}$. Thereafter, the algorithm optimally chooses a step $\tau_{t}$ towards $s_{t}$ from $x_{t-1}$ and sets $x_{t}=x_{t-1}+\tau_{t}(s_{t}-x_{t-1})$ for $\tau_{t}\in[0,1]$. The advantage of this method over other first-order methods like gradient descent is that it avoids projection (which can often be expensive).
1:An initial feasible point in the dual space $y^{0}=(\beta^{0},p^{0})\in\mathbb{R}^{n+m}_{\geq 0}$ such that $\sum_{j}p^{0}_{j}=\sum_{i}B_{i}$
2:for $t=1,2,\ldots$ do
3:     $y^{t+1}=\arg\max_{y\in Y}\,\langle\nabla f(y^{t}),y-y^{t}\rangle$
4:     if $\langle\nabla f(y^{t}),y^{t+1}-y^{t}\rangle=0$ then return $y^{t}$
5:     end if
6:end for
Algorithm 1 Greedy Frank-Wolfe (GFW) algorithm
For convex minimization, the FW method has been immensely popular, for example in machine learning application (Jaggi 2013), and it has previously been used to compute CE in various settings with goods (Gao and Kroer 2020, Panageas et al. 2021). In the goods setting, one must be careful about stepsize selection, in order to guarantee convergence; typically a step is taken from the current iterate in the direction of the point maximizing (or minimizing in the convex case) the linearization. Interestingly, in the chores setting this is not the case; because we are maximizing a convex objective, we can greedily move directly to the point returned by the linear maximization. In particular, for our (Chores Dual Redundant) program, we can greedily choose our current iterate point to be the maximizer $s_{t}$ of the linear approximation of the function $f$ around $y_{t-1}$. That is, we set $y_{t}=s_{t}$ where $s_{t}={\operatorname{\mathrm{arg\,max}}}_{s\in Y}\langle\nabla f(y_{t-1}),s-y_{t-1}\rangle$. This yields 6. We call this algorithm the Greedy Frank Wolfe (GFW) algorithm.
GFW has previously been studied for generic problems of maximizing a convex function in the context of certain machine learning applications (Mangasarian 1996, Journée et al. 2010). It is instructive to briefly understand why we can be greedy with this method. Because we have a convex function, the following inequality holds,
$f(y^{t+1})-f(y^{t})\geq\langle\nabla f(y^{t}),y^{t+1}-y^{t}\rangle.$ (9)
Note that the right-hand side is exactly the objective in the line 2 of 6, and thus, since $y^{t}$ is a feasible solution, the RHS is weakly greater than zero. Moreover, the solution strictly improves, unless we have already found a local maximum (thus a KKT point and a CE). This immediately implies that we have a monotonically-improving sequence of iterates. Secondly, note that there always exists a vertex of $Y$ which is optimal, and since there is a finite set of vertices in $Y$, this implies that the method converges in finite time. These observations were formalized by Mangasarian (1996):
Theorem 4 (Mangasarian (1996)).
Let $f$ be a differentiable convex function that is bounded above on a polyhedron $Y$. Then, 6 generates a sequence of iterates $\{y^{t}\}_{t\geq 1}$ such that
  • (i)
    $\{y^{t}\}_{t\geq 1}$ is well-defined;
  • (ii)
    $f(y^{t})$ strictly increases as $t$ increases;
  • (iv)
    The algorithm terminates with a $y^{T}$ satisfying the first-order optimality condition
    $\langle\nabla f(y^{T}),y-y^{T}\rangle\leq 0\quad\text{for all }y\in Y$
    in a finite number of iterations $T$.
Figure 2: The trajectory of GFW until termination on a $2$-agent-$8$-chore instance, starting from a feasible interior point. The gray region is the feasible set projected onto the $(\beta_{1},\beta_{2})$ subspace. The blue lines with arrows denote the trajectory of GFW. The orange arrows and dotted lines denote gradient ascending directions of $f$ and corresponding planes perpendicular to these directions, respectively.
In Fig. 2, we illustrate a trajectory of the GFW algorithm on an instance with 2 agents and 8 chores. The earning requirements of agents are set to be one and disutilities are drawn from the [0, 1] uniform distribution.
While the above guarantees finite-time convergence, it does not give a rate at which this convergence occurs. Later, some convergence-rate results were shown by Journée et al. (2010). Most related to our chores CE setting, they show that for strongly convex functions, GFW has a best-iterate convergence rate of $O(1/(\sigma_{f}\epsilon^{2}))$, where the convergence is in terms of some iterate $k\in\{1,\ldots,T\}$ having a small stepsize $\|y^{k+1}-y^{k}\|\leq\epsilon$, and $\sigma_{f}$ is the strong convexity parameter. However, our objective is not strongly convex, and thus we cannot apply the results from Journée et al. (2010) to the chores CE setting. In the next sections, we will show how to generalize the result of Journée et al. (2010) to a class of Bregman-strongly-convex functions. We then show that our objective indeed satisfies this generalized notion of strong convexity, and thus is guaranteed a small multiplicative stepsize difference. We go on to show that this implies convergence to chores CE. In the appendix, we also give an alternative proof of this convergence directly from arguing about CE structure, without going through Bregman strong convexity.
4.1 General GFW Convergence Results for Convex Maximization
In this part, we generalize our Chores EG dual problem and show general convergence results for $\log$-type convex maximization over any nonempty closed convex set. We show that the multiplicative difference between two consecutive iterates is guaranteed to be smaller than any $\epsilon>0$ after $O(\frac{1}{\epsilon^{2}})$ number of iterations. We then show in the following section that such multiplicative “stepsize convergence” implies convergence to an approximate chores CE.
Formally, we consider the following $\log$-type convex maximization problem:
$\max_{x\in\mathbb{R}^{n}_{>0}}$ $g(x):=-\sum_{i=1}^{n}a_{i}\log{x_{i}}$ (10)
s.t. $x\in\mathcal{X},$
where $a_{i}>0,\;\forall\;i\in[n]$ and $\mathcal{X}$ represents any nonempty closed convex set. We only assume a finite upper bound for $g(x)$, i.e., there exists a $g^{*}<+\infty$ such that $g(x)\leq g^{*}$ for all $x\in\mathcal{X}$.
To do this, we first generalize the results from Journée et al. (2010) on convergence for strongly convex objectives to a convergence result for objectives that are strongly convex relative to some strictly convex differentiable function $h$. This result is quite general, and may be of interest outside of our CE setting.
Theorem 5.
Let $f:\mathcal{X}\rightarrow\mathbb{R}$ be a continuously differentiable function that is $\alpha$-strongly convex relative to $h$, where $\mathcal{X}$ is a nonempty closed convex set, and bounded above by $f^{*}<+\infty$ on $\mathcal{X}$. If we maximize $f$ over $\mathcal{X}$ using the GFW algorithm starting from an arbitrary feasible point $x^{0}$, then for any $\epsilon>0$, we can find a pair of consecutive iterates $(x^{t},x^{t+1})$ such that $D_{h}(x^{t+1}\,\|\,x^{t})\leq\epsilon$ after at most $\left\lceil\frac{f^{*}-f(x^{0})}{\alpha\epsilon}\right\rceil$ iterations.
Proof.
We first prove that the GFW is well-defined if $f$ is convex and bounded above by $f^{*}<+\infty$. Since we have for each iteration $t$
$0\leq\langle\nabla f(x^{t}),x^{t+1}-x^{t}\rangle\leq f(x^{t+1})-f(x^{t})\leq f^{*}-f(x^{t})<+\infty,$
$\mathcal{X}$ has to be bounded on the direction of $x^{t+1}-x^{t}$. Also, $x^{t+1}$ is attainable since $\mathcal{X}$ is closed.
Let $T=\left\lceil\frac{f^{*}-f(x^{0})}{\alpha\epsilon}\right\rceil$. Since $x^{t+1}$ maximizes $\langle\nabla f(x^{t}),x\rangle$ over $\mathcal{X}$ and $x^{t}\in\mathcal{X}$,
$\langle\nabla f(x^{t}),x^{t+1}-x^{t}\rangle\geq 0\quad\forall t=0,1,\ldots,T.$
By $\alpha$-strong convexity with respect to $D_{h}(\cdot\,\|\,\cdot)$, we have
$f(x^{t+1})\geq f(x^{t})+\langle\nabla f(x^{t}),x^{t+1}-x^{t}\rangle+\alpha D_{h}(x^{t+1}\,\|\,x^{t})\quad\forall\;t=0,1,\ldots,T.$
Summing Section 4.1 up over $t=0,\ldots,T$ and applying Section 4.1, we have
$\alpha T\min_{0\leq t\leq T}D_{h}(x^{t+1}\,\|\,x^{t})\leq\alpha\sum_{t=0}^{T}D_{h}(x^{t+1}\,\|\,x^{t})\leq f(x^{T+1})-f(x^{0})\leq f^{*}-f(x^{0}).$
Thus, there exists a $t\leq T$ such that $D_{h}(x^{t+1}\,\|\,x^{t})\leq\epsilon$ because $\min_{0\leq t\leq T}D_{h}(x^{t+1}\,\|\,x^{t})\leq\frac{f^{*}-f(x^{0})}{\alpha T}\leq\epsilon$. ∎
Now we consider the $\log$-type convex maximization case in Eq. 10. Let the reference function be $h(x)=-\sum_{i}\log{x_{i}}$. The Bregman divergence associated with $h$ is
$D_{IS}(y\,\|\,x)=D_{h}(y\,\|\,x)=\sum_{i}\left(-\log{\frac{y_{i}}{x_{i}}}+\frac{y_{i}}{x_{i}}-1\right),$ (11)
where $D_{IS}(\cdot\|\,\cdot)$ refers to the Itakura–Saito distance. Then, for the $\log$-type function in Eq. 10, we trivially have
$g(y)\geq g(x)+\langle\nabla g(x),y-x\rangle+\min_{i}a_{i}\;D_{IS}(y\|x),$ (12)
since $g(y)-g(x)-\langle\nabla g(x),y-x\rangle$ can be seen as a sum over component-wise Itakura-Saito distances, each weighted by $a_{i}$, and due to nonnegativity, taking the minimum over $a_{i}$ can only make each term smaller.
Now, we note that $D_{IS}(y\,\|\,x)$ is lower bounded by $\frac{1}{3}\sum_{i}\big{(}\frac{y_{i}}{x_{i}}-1\big{)}^{2}$ when $\frac{y_{i}}{x_{i}}$ is close to $1$ for each $i$ (or, equivalently, $D_{IS}(y,\|\,x)$ is close to 00). In particular, we have
$\frac{1}{3}\sum_{i}\left(\frac{y_{i}}{x_{i}}-1\right)^{2}\leq D_{IS}(y\,\|\,x)$ (13)
if $-\log{\frac{y_{i}}{x_{i}}}+\frac{y_{i}}{x_{i}}-1\leq c:=\sqrt{3}-1-\frac{1}{2}\log{3}$ for all $i$.
With this, we obtain the following result.
Theorem 6.
Suppose we solve Eq. 10 using the GFW algorithm starting from an arbitrary feasible point $x^{0}$. Assume that there exists a finite upper bound $g^{*}<+\infty$. Then, for any $\epsilon>0$, we can find a pair of iterates $(x^{t},x^{t+1})$ such that $\sum_{i}\big{(}\frac{x^{t+1}_{i}}{x^{t}_{i}}-1\big{)}^{2}\leq\epsilon^{2}$ after at most
$\left\lceil\frac{3(g^{*}-g(x^{0}))}{\min_{i}a_{i}}\frac{1}{\epsilon^{2}}+\frac{g^{*}-g(x^{0})}{c\min_{i}a_{i}}\right\rceil$
iterations, where $c=\sqrt{3}-1-\frac{1}{2}\log{3}>0$.
Proof.
Let $T=\left\lceil\frac{3(g^{*}-g(x^{0}))}{\min_{i}a_{i}}\frac{1}{\epsilon^{2}}+\frac{g^{*}-g(x^{0})}{c\min_{i}a_{i}}\right\rceil$. We prove this result for two separate cases: $\epsilon>\sqrt{3c}$ and $\epsilon\leq\sqrt{3c}$.
If $\epsilon>\sqrt{3c}$, we can find a pair of $(x^{t},x^{t+1})$ such that $D_{IS}(x^{t+1}\|x^{t})\leq c<\epsilon^{2}/3$ after at most $\left\lceil\frac{g^{*}-g(x^{0})}{c\alpha}\right\rceil\leq T$ iterations by Theorem 5.
Since $-\log{a}+a-1\geq 0$ for all $a>0$, we have $-\log{\frac{x^{t+1}_{i}}{x^{t}_{i}}}+\frac{x^{t+1}_{i}}{x^{t}_{i}}-1\leq c$ for all $i\in[n]$. Hence, by Eq. 13,
$\frac{1}{3}\sum_{i}\left(\frac{x^{t+1}_{i}}{x^{t}_{i}}-1\right)^{2}\leq D_{IS}(x^{t+1}\,\|\,x^{t})\leq\frac{\epsilon^{2}}{3}$
If $\epsilon\leq\sqrt{3c}$, then $D_{IS}(x^{t+1}\,\|\,x^{t})\leq\frac{\epsilon^{2}}{3}\leq c$ after $\frac{3(g^{*}-g(x^{0}))}{\min_{i}a_{i}}\frac{1}{\epsilon^{2}}\leq T$ by Theorem 5. Again, we have $-\log{\frac{x^{t+1}_{i}}{x^{t}_{i}}}+\frac{x^{t+1}_{i}}{x^{t}_{i}}-1\leq c$ for all $i\in[n]$, and by Eq. 13, $\frac{1}{3}\sum_{i}\left(\frac{x^{t+1}_{i}}{x^{t}_{i}}-1\right)^{2}\leq D_{IS}(x^{t+1}\,\|\,x^{t})\leq\frac{\epsilon^{2}}{3}$. ∎
Remark.
The result from Journée et al. (2010) does not apply to our setting since our objective is not strongly convex. It is possible that the extrapolation idea from Gao and Kroer (2020) could be applied in our setting in order to get some form of strong convexity and apply the results from Journée et al. (2010). However, even if this approach were to work, it would lead to significantly worse polynomial dependence on $n$.
4.2 Convergence to Approximate Competitive Equilibrium
With Theorem 6 in hand, we are ready to show polynomial-time convergence to approximate competitive equilibrium. We attain this by showing that an $\epsilon$-small multiplicative stepsize implies that we have reached an $\epsilon$-approximate competitive equilibrium.
Before showing this implication, first we show that prices are always strictly positive.
Claim 1.
The prices at a given iteration $t$ satisfy $p^{t}_{j}>0$ for all $j\in[m]$.
Proof.
We prove this by contradiction. Assume there exists a $j\in[m]$, such that $p^{t}_{j}=0$. Then note that $p^{t}_{j}/d_{ij}<\beta^{t}_{i}$ for all $i\in[n]$ (this follows from the fact that every $\beta^{t}_{i}$ is strictly positive). Due to the strict inequality, there must exist some (possibly small) $\gamma>0$ such that if we increase $p_{j}^{t}$ by $\gamma$, and decrease the price of every other good by $\gamma/(\sum_{i}B_{i}-1)$, then we still have a feasible set of prices $p^{\prime}$, and $\max_{j\in[m]}p^{t}_{j}/d_{ij}$ strictly decreases for all $i\in[n]$. Let $\beta^{\prime}_{i}=\max_{j\in[m]}p_{j}^{\prime}/d_{ij}$. Note that $\beta^{\prime}_{i}<\beta^{t}_{i}$ for all $i\in[n]$ and thus $\sum_{i\in[n]}B_{i}\cdot\beta^{\prime}_{i}/\beta^{t-1}_{i}<\sum_{i\in[n]}B_{i}\cdot\beta^{t}_{i}/\beta^{t-1}_{i}$ which contradicts the optimality of $\beta^{t}$ when solving the GFW LP. ∎
Lemma 2.
Let $\{(p^{t},\beta^{t})\}_{t\geq 0}$ be a sequence of iterates generated by the GFW algorithm. Let $x^{t}$ be the optimal dual variable of the LP subproblem at the $t$-th iteration and
$\bar{x}^{t}=\frac{\sum_{i}B_{i}}{\sum_{i}B_{i}\frac{\beta^{t}_{i}}{\beta^{t-1}_{i}}}x^{t}.$
If $\big{\lvert}\frac{\beta^{t}_{i}}{\beta^{t-1}_{i}}-1\big{\rvert}\leq\epsilon$ for all $i\in[n]$, then $(p^{t},\bar{x}^{t})$ form an $\epsilon$-strongly approximate CE.
Proof.
Given the $t$-th iterate $(p^{t-1},\beta^{t-1})$, the next iterate of the GFW algorithm is generated by the solution to the following LP:
$\min_{p,\beta\geq 0}$ $\sum_{i}B_{i}\frac{\beta_{i}}{\beta^{t-1}_{i}}$ (14)
s.t. $\quad\;\;p_{j}\leq d_{ij}\beta_{i},\quad\forall\,i,j,$ $[x_{ij}]$
$\sum_{j}p_{j}=\sum_{i}B_{i}.$ $[\mu]$
Let $x$ and $\mu$ denote the dual variables corresponding to the first and second constraints of Eq. 14, respectively. We use $(p^{t},\beta^{t})$ and $(x^{t},\mu^{t})$ to denote the optimal primal and dual variables at the $t$-th iteration, respectively.
By KKT conditions and strong duality of LP, $(p,\beta)$ is an optimal solution to Eq. 14 if and only if there exists a pair of $(x,\mu)$ such that $(p,\beta,x,\mu)$ satisfies
  1. (i)
    $p,\beta,x\geq 0$;
  2. (ii)
    $p_{j}\leq d_{ij}\beta_{i}$ and $x_{ij}\left(p_{j}-d_{ij}\beta_{i}\right)=0$ for all $i\in[n],j\in[m]$;
  3. (iii)
    $\sum_{j}p_{j}=\sum_{i}B_{i}$ and $\mu\left(\sum_{j}p_{j}-\sum_{i}B_{i}\right)=0$;
  4. (iv)
    $\sum_{i}x_{ij}+\mu\geq 0$ and $p_{j}\left(\sum_{i}x_{ij}+\mu\right)=0$ for all $j\in[m]$;
  5. (v)
    $\frac{B_{i}}{\beta^{t-1}_{i}}-\sum_{j}d_{ij}x_{ij}\geq 0$ and $\beta_{i}\big{(}\frac{B_{i}}{\beta^{t-1}_{i}}-\sum_{j}d_{ij}x_{ij}\big{)}=0$ for all $i\in[n]$.
Recall that $(p^{t},\beta^{t})$ is an optimal solution to Eq. 14. By condition (ii), we have $\frac{p^{t}_{j}}{d_{ij}}=\beta^{t}_{i}\geq\frac{p^{t}_{j^{\prime}}}{d_{ij^{\prime}}}$ for all $j^{\prime}\in[m]$ if $x^{t}_{ij}>0$ for all $i,j$. This guarantees every buyer is allocated her optimal bundle subject to the amount of budget she earns.
Second, we have
[…]
which leads to $(1-\epsilon)B_{i}\leq\sum_{j}p_{j}x_{ij}\leq(1+\epsilon)B_{i}\leq\frac{B_{i}}{1-\epsilon}$ since $\big{|}\frac{\beta^{t}_{i}}{\beta^{t-1}_{i}}-1\big{|}\leq\epsilon$ for all $i\in[n]$. That is, every buyer approximately earns her budget.
Next, by summing up the equality in condition (iv), we obtain
$\sum_{i}\sum_{j}p^{t}_{j}x^{t}_{ij}+\mu^{t}\sum_{j}p^{t}_{j}=0.$ (16)
Combining this with Eq. 15 and condition (iii), we have $\mu^{t}=-{\sum_{i}B_{i}\frac{\beta_{i}}{\beta^{t-1}_{i}}}\Big{/}{\sum_{i}B_{i}}$ and
$\sum_{i}x^{t}_{ij}=-\mu=\frac{\sum_{i}B_{i}\frac{\beta_{i}}{\beta^{t-1}_{i}}}{\sum_{i}B_{i}}\quad\text{if }p^{t}_{j}>0.$ (17)
By 1, this holds for all $j\in[m]$. It follows that
$\sum_{i}\bar{x}^{t}_{i}=\frac{\sum_{i}B_{i}}{\sum_{i}B_{i}\frac{\beta^{t}_{i}}{\beta^{t-1}_{i}}}\sum_{i}x^{t}_{i}=1\quad\mbox{for all }j\in[m].$ (18)
This means $\bar{x}$ is a feasible allocation (exactly allocating all chores).
Therefore, the lemma follows by the definition of $\epsilon$-strongly approximate CE. ∎
Therefore, by Lemmas 2 and 6, we can conclude that the GFW algorithm finds an $\epsilon$-approximate CEEI in $\left\lceil\frac{3(g^{*}-g(x^{0}))}{\min_{i}a_{i}}\frac{1}{\epsilon^{2}}+\frac{g^{*}-g(x^{0})}{c\min_{i}a_{i}}\right\rceil$ iterations. Observe that $g^{*}-g(x^{0})$ can be upper bounded by
$\sup_{p,\beta\in Y}-\sum_{i}B_{i}\log(\beta_{i}).$
To this end, note that for each agent $i$, $1/\beta_{i}\leq\min_{j}d_{ij}/p_{j}$ (implied by the constraints of our program). Since $\sum_{j}p_{j}=\sum_{i}B_{i}$, this implies that there exists at least one chore $j$, such that $p_{j}\geq\sum_{i}B_{i}/m$, implying that $d_{ij}/p_{j}\leq mD/(\sum_{i}(B_{i}))$, where $D=\max_{i,j}d_{ij}$. Therefore, we have
$-\sum_{i}B_{i}\log(\beta_{i})=\sum_{i}B_{i}\log(1/\beta_{i})\leq\sum_{i}B_{i}\log(mD/(\sum_{i}B_{i}))\leq n\max_{i}B_{i}\log(mD/\min_{i}B_{i}).$
Theorem 7 (Complexity of finding an approximate CE).
The GFW algorithm finds an $\epsilon$-strongly approximate CE in $\mathcal{O}(n\frac{\max_{i}B_{i}}{\min_{i}B_{i}}\log(\frac{mD}{\min_{i}B_{i}})/\epsilon^{2})$ iterations.
For CEEI, our results (Theorem 7) improve the current state-of-the-art convergence rate in terms of the number of iterations relative to $\epsilon$: The EPM by Boodaghians et al. (2022) requires $\mathcal{\tilde{O}}(n^{3}/\epsilon^{2})$ iterations, and the combinatorial algorithm by Chaudhury et al. (2022) requires $\mathcal{\tilde{O}}(nm/\epsilon^{2})$ iterations, in contrast to $\mathcal{{\tilde{O}}}(n/\epsilon^{2})$ iterations of GFW [^5] For CEEI, note that $\max_{i}B_{i}=\min_{i}B_{i}=1$., to reach an approximate CE. Furthermore, we expect the cost of each iteration to be cheaper for GFW. In GFW, we only need to solve an LP in every iteration. Moreover, the approximate solution of the LP poses no numerical issues. In contrast, solving the convex QP in EPM is a more challenging task. It is hard to find exact running times for general convex QPs, and even quite good approximate solutions can lead to serious numerical issues, as we show in Section 5. Further, we believe that the LP solved in every iteration of GFW can be solved (in worst case) using algorithms faster than the ones used for general-purpose LPs van den Brand (2020), and we leave this as an interesting question for future research.
The result in Theorem 7 can also be recovered with mildly stronger guarantees through a primal-dual analysis of the GFW (See Remark Remark). The analysis in Remark Remark gives us further insights on each run of GFW: for instance, in each iteration of GFW where we do not yet have a good approximate CE point, we make a large improvement on the objective function (5).
5 Experiments
In Section 4, we showed that the GFW algorithm guarantees convergence to a strongly approximate CE in polynomial time, and converges to an exact CE in finite number of iterations, while requiring solving only one LP per iteration. In contrast, the prior numerical state-of-the-art method, the EPM of Boodaghians et al. (2022) requires solving a convex quadratic program (QP) at every iteration.
In this section, we investigate the numerical performance of GFW as compared to the EPM, when both utilize state-of-the-art optimization software for solving LPs and QPs (in particular, using Gurobi). We remark that we do not provide comparisons with the combinatorial algorithm (Chaudhury et al. 2022). This is primarily attributed to the numerical issues that arise with the arithmetic performed by combinatorial algorithms–in particular, the bit-length of the prices and the allocation can grow exponentially with the iterations in the combinatorial algorithms, as conjectured by Duan and Mehlhorn (2015). [^6] In private communication, we are aware that Omar Darwish and Kurt Mehlhorn have an implementation of the combinatorial algorithm for Arrow-Debreu markets with goods, and they confirm the conjecture in Duan and Mehlhorn (2015).
In 10 we give pseudocode for the EPM for computing an exact CE for chores. The set $\mathcal{D}^{+}$ in the pseudocode denotes the set of all possible disutility profiles corresponding to any feasible allocation or over-allocation, i.e.,
$\mathcal{D}^{+}=\Big{\{}d\in\mathbb{R}^{n}_{\geq 0}\;|\;\exists\;x\in\mathbb{R}^{n\times m}_{\geq 0}\mbox{ such that }d_{i}\geq\sum\nolimits_{j}d_{ij}x_{ij}\;\forall\;i\in[n],\sum\nolimits_{i}x_{ij}=1\;\forall\;j\in[m]\Big{\}}.$
[…]
2:for $k=1,2,\ldots$ do
3:     if $d^{k}$ is infeasible then
4:         $d^{k,*}={\arg\min}_{d\in\mathcal{D}^{+}\cap\{d|d\geq d^{k}\}}\big{\lVert}d-d^{k}\big{\rVert}$ 77footnotemark: 7
5:         $a^{k}\leftarrow{(d^{k,*}-d^{k})n}\big{/}{\langle d^{k,*}-d^{k},d^{k,*}\rangle}$
6:         $d^{k+1}\leftarrow{1}\big{/}{a^{k}}$
7:     else
8:         return $(d^{k},a^{k-1})$
9:     end if
10:end for
Algorithm 2 Exterior Point Method (EPM)
Experimental Setup
We compare the algorithms on a variety of instances. First, we consider randomly generated instances. For a given instance size $n$, we construct $n$ buyers, each with a budget of $1$. Then, we generate $n$ chores. This means that disutilities are an $n\times n$ matrix $D$. Each cell $d_{ij}$ in $D$ is sampled i.i.d. from some distribution $F$: $d_{ij}\sim F$. We consider five different distributions of disutilities: (1) uniform on $[0,1)$; (2) log-normal associated with the standard normal distribution $\mathcal{N}(0,1)$; (3) truncated normal associated with $\mathcal{N}(0,1)$ and truncated at ${10}^{-3}$ and $10$ standard deviations from 00; (4) exponential with the scale parameter $1$; (5) uniform random integers on $\{1,\ldots,1000\}$. For each distribution, we consider instance sizes $n\in\{2,50,100,150,200,250,300\}$. For each choice of distribution and $n$, we generate 100 instances.
Secondly, we generate a set of chores instances by considering a dataset of bids from PC members at the AAMAS conference on potential papers to review, obtained from PrefLib (Mattei and Walsh 2013). We use the 2021 dataset, which has 596 PC members included in the dataset. Each PC member scored each paper ordinally as {yes >>> maybe >>> no response >>> no >>> conflict}. In order to generate real-valued disutilities, we convert these into the values {1, 3, 5, 7, $4000$}. We then generate two different datasets: In the first, we use these 5 values directly, and sample subsets of $n$ PC members and $n$ papers as follows: first uniformly choose one paper at random, then choose a subset of papers (of size $n$) which are the most similar to it (considering the PC members’ responses as feature vectors of the paper), finally choose the subset (of size $n$) of PC members that have the largest number of non-conflict responses on the chosen papers. In the second, we additionally add Gaussian noise to each valuation, using $\mathcal{N}(0,0.04)$. For both datasets, we consider instance size $n\in\{2,50,100,150,200,250,300\}$. For each dataset and $n$, we generate 100 instances.
We ran all experiments on a personal desktop that uses the Apple M1 Chip and has 8GB RAM. We used Gurobi version 11.0.0. Since the tolerance parameters for Gurobi models may influence the accuracy of the solutions per iteration and thus the performance of EPM, we tuned those parameters. We considered 4 parameters: we tuned FeasibilityTol in the range of $10^{-9}$ to $10^{-6}$, OptimalityTol in the range of $10^{-9}$ to $10^{-6}$, BarConvTol in the range of 00 to $10^{-8}$, and BarCorrectors in the range of 00 to $10^{4}$  [^8] The default setting chooses this parameter automatically, depending on problem characteristics. . For our final experiments, we set these parameters as FeasibilityTol = $10^{-6}$ (default), OptimalityTol = $10^{-6}$ (default), BarConvTol = 00 (minimum), and BarCorrectors as default. This is because shrinking BarConvTol significantly increases the accuracy of solutions; but decreasing the other tolerance parameters, or increasing BarCorrectors, greatly increases the time Gurobi spends per iteration, thus harming the EPM performance, with no significant improvement to the solution accuracy. See Appendix C for plots illustrating this point.
In our experiments, we measure the computational cost required to find a strongly-approximate CE and an exact CE. In particular, we measure the number of iterations and the running time. Recall the definitions of $\epsilon$-approximate CE and exact CE in Section 2. We consider any condition with $\epsilon\leq 10^{-6}$ as a condition that holds exactly. We consider any CE with the first condition relaxed to $\epsilon\leq 0.01$ (and the other two conditions exactly satisfied) as a strongly-approximate CE. To measure the equilibrium approximation quality $\epsilon$ at a given iteration, we compute a pair $(p,x)$ corresponding to the current iterate for each algorithm. For GFW, we compute $(p,x)$ based on Lemma 2, where $x$ is given by the optimal dual variable of the corresponding LP. For EPM, we solve an LP to find a feasible $x$ corresponding to the final $d^{k}$ if the algorithm finds an exact CE, and retrieve $x$ corresponding to $d^{k,*}$ if $(d^{k,*},a^{k})$ corresponds to a strongly-approximate CE. We compute prices $p$ for EPM via $p_{j}=\min_{i}a_{i}d_{ij}$ (Boodaghians et al. 2022).
Though the EPM is supposed to reach an exact CE when it finds a feasible (or over-allocated) disutility profile, we found in our experiments that an unignorable numerical error can be caused by the approximate solution of QPs. If the projection step in 10 cannot be solved accurately, this can cause $a^{k-1}$, the normal vector of the generated hyperplane, to point in the wrong direction. This leads to an inaccurate price vector $p^{*}$. Moreover, let $d^{k}$ be the true iterate with a perfect-precision QP solution, and let $\tilde{d}^{k}$ be the “solution” found by Gurobi at a problematic iteration. The Gurobi solution $\tilde{d}^{k}$ may end up being feasible even if $d^{k}$ is not, or vice versa. This may lead to incorrect termination decisions in either case (in which case our attempt to compute a corresponding CE allocation fails). This is a critical issue, because we have no way to proceed once this happens. Considering this issue, we say that the EPM fails if we cannot find a corresponding strongly approximate CE even though the EPM ostensibly terminated with one. In our experiments, we also measure the fraction of instances successfully solved without this issue.
Experiment Results and Discussion
Our results on the five different distributions for i.i.d. random disutilities are shown in Fig. 3.
uniform(0,1) Refer to caption Refer to caption Refer to caption lognormal Refer to caption Refer to caption Refer to caption truncated normal Refer to caption Refer to caption Refer to caption exponential Refer to caption Refer to caption Refer to caption randint(1,1000) Refer to caption Refer to caption Refer to caption
Figure 3: Numerical results on iid randomly generated disutility matrix instances. Each row corresponds to a different disutility distribution (noted on the left). Left: Average number of iterations to solve a given instance size. Middle: Average wall clock running time to solve a given instance size (using Gurobi for both algorithms). Right: Fraction of instances solved for each algorithm. GFW solves every instance for every distribution.
In the left column of plots, we show the average number of iterations taken for each algorithm before finding a solution. In the middle column of plots, we show the mean running time. For both the left and middle column of plots, we only average over instances that were solved by EPM (and if EPM reaches the largest number on the y axis it denotes that no instances were solved at a given size). The right column shows the fraction of instances solved for each instance size.
In terms of the number of iterations taken by each algorithm, we see that that they are very comparable: in all cases where both algorithms solve some instances, the number of iterations is quite low: always below 30 iterations. When it does work, the EPM seems to take slightly fewer iterations than GFW in some cases. In terms of runtime, GFW is generally faster than EPM: in every case where EPM solves at least one instance, we have the GFW is faster, sometimes by a factor of 2-4. By far the most important result in terms of practicality is the right column of plots. We see that GFW successfully solves every instance we tried, out of the 3500 instances generated in total. In contrast, EPM fails on a significant number of instances, sometimes failing on every instance for larger instance sizes, when attempting to compute exact CE.
The results on our semi-synthetic AAMAS bidding instances are shown in Fig. 4. Again we see that the number of iterations to reach CE is quite comparable across GFW and EPM, on the instances where EPM succeeds. The running time results are also similar: both for exact and approximate CE, GFW is faster than EPM in both the original data setting, and the added-noise setting. Notably, both algorithms are very fast, terminating in less than ten seconds for the original data, and GFW still terminates in less than 15 seconds for the added-noise setting. On the original data, both algorithms solve every instance. This is likely due to the fact that the original data is a very special type of instance: there are only 5 possible distinct disutility values, and every buyer has a lot of chores with value 1. Thus, the equilibrium is very simple, which is also evident in the fact that the mean number of iterations is 3.5 even for the largest instances. On the added-noise data, we see again that EPM exhibits significant numerical issues, failing on a large fraction of instances, especially for exact CE.
Overall, we conclude that GFW is the first highly practical algorithm: it solves every instance in very little time and in few iterations, even for instances where the size of the disutility matrix is $300\times 300$. In comparison, EPM fails to solve many instances in our datasets. As discussed before, we do caution that the success of EPM is highly dependent on the performance of the Gurobi QP solver, and in particular its accuracy. We found that the QP performance of Gurobi can vary substantially across different machines, and unpredictably so. Due to this so-called “performance variability”, it is difficult to conclusively evaluate the wall-clock performance of the algorithms, as both algorithms rely on the performance of Gurobi on large-size instances. In particular, some preliminary experiments suggests that the wall-clock performance of EPM improves relative to the wall-clock performance of GFW if one adds more RAM. However, we found that even in this case, EPM still has a significant failure rate, since Gurobi still outputs approximate QP solutions, and this still leads to numerical errors that break EPM.
original Refer to caption with noise Refer to caption
Figure 4: Numerical results on AAMAS PC member bidding data. The top row uses the “original” valuations. The bottom row adds Gaussian noise. Left: Average number of iterations to solve a given instance size. Middle: Average wall clock running time to solve a given instance size. Right: Fraction of instances solved.
6 Conclusion and Discussion
We develop a new convex maximization program for computing CE in Fisher markets with chores, and showed that it satisfies some dual-like properties with the concave minimization program for computing CE. We then derived a new redundant constraint which allows us to avoid directions that lead to infinitely positive objective. This yielded the first mathematical program for finding CE for chores such that it is possible to run iterative optimization methods over a convex feasible region. We then introduced the GFW method for computing CE. From prior results on convex maximization over polyhedra, it follows that this method terminates in a finite number of iterations. We then showed that, in fact, the method finds an $\epsilon$-approximate CE in $O(\frac{1}{\epsilon^{2}})$ iterations, while requiring solving only a simple LP at every iteration. We also gave a more general convergence result that generalizes prior results for strongly convex functions to Bregman-strongly-convex functions. Finally, we showed through extensive numerical experiments that our method is highly practical: it solved every instance that we tried in a minute or less and is both robust and simple to implement. In contrast, we showed that the prior state-of-the-art method, the EPM, often fails to solve instances altogether, due to numerical accuracy problems that break the algorithm.
Challenge of moving from approximate CE to exact CE.
The obvious open problem is to settle the exact complexity of finding a CE in the chores market. A natural attempt would be to find a good approximate CE and then argue how to move to a “nearby” exact CE. Unfortunately, such approaches can face the issue that there are instances where an approximate CE can be very far from an exact CE (See Appendix B for an example).
Therefore, one may need to search for additional properties while looking out for a good approximate CE/ CEEI. Note that our GFW method indeed finds an exact CE in the chores market in finite time. In fact, it finds an exact CE in very few iterations in our experiments. Thus, it would be interesting to investigate instances where the time taken by GFW to reach an exact CE is exponential. We believe that this can improve our understanding of “hard instances” (if any) for iterative methods for computing a CE.
Despite the above hurdle, we believe that finding a $\epsilon$-approximate CE in the chores market with a better time complexity dependence on $\frac{1}{\epsilon}$ is of theoretical interest. An ideal goal would be to have a poly-logarithmic dependence on $\frac{1}{\epsilon}$, but any polynomial improvement would be a stepping stone (currently, all methods have a dependence of $\mathcal{O}(\frac{1}{\epsilon^{2}})$).
Acknowledgement
Christian Kroer was supported by the Office of Naval Research awards N00014-22-1-2530 and N00014-23-1-2374, and the National Science Foundation awards IIS-2147361 and IIS-2238960. Ruta Mehta was supported by the NSF grant CCF-2334461.
Appendix A A Primal-Dual Analysis of the Convergence of the Frank Wolfe Method to an Approximate Equilibrium
Since $f$ does not depend on $p$, we ease notation in this by section by writing $f(\beta)$ for the (Chores Dual) objective function, which was previously defined as depending on both $p$ and $\beta$.
Let $\{(p^{t},\beta^{t})\}_{t\geq 0}$ be the sequence of iterates generated by the greedy Frank Wolfe algorithm. Then, given $\beta^{t-1}_{i}$ for all $i\in[n]$, iteration $t$ of the Frank Wolfe algorithm attempts to find the vector $\beta^{t}$ such that
$\min_{\beta}$ $\sum_{i\in[n]}B_{i}\cdot\frac{\beta_{i}}{\beta^{t-1}_{i}}$ (19)
s.t. $p_{j}-\beta_{i}d_{ij}\leq 0$ $\forall i,j$
$\quad\sum_{j\in[m]}p_{j}=\sum_{i\in[n]}B_{i}=B$
$\quad p_{j}\geq 0$ $\forall j$
$\quad\beta_{i}\geq 0$ $\forall i$
Let $x^{t}_{ij}$ be the optimal dual variable corresponding to the constraint $p_{j}-\beta_{i}d_{ij}\leq 0$, and $\delta^{t}$ be the optimal dual variable corresponding to the constraint $\sum_{j\in[m]}p_{j}=1$ in the $t$-th iteration of the greedy Frank Wolfe algorithm.
Claim 2.
For all $i\in[n]$, we have $\sum_{j\in[m]}d_{ij}x^{t}_{ij}=B_{i}/\beta^{t-1}_{i}$.
Proof.
Under stationarity conditions of the optimum solutions, we have for all $i\in[n]$, $B_{i}/\beta^{t-1}_{i}=\sum_{j\in[m]}d_{ij}x^{t}_{ij}$, whenever $\beta^{t}_{i}>0$. Thus, it suffices to show that $\beta^{t}_{i}>0$ for all $i\in[n]$. Assume otherwise, and $\beta^{t}_{\ell}=0$. Since $\beta^{t}_{\ell}=\max_{j\in[m]}p^{t}_{j}/d_{\ell j}$, this implies that $p^{t}_{j}=0$ for all $j\in[m]$, implying $\sum_{j\in[m]}p^{t}_{j}=0$, which is a contradiction. ∎
We now make another technical observation.
Claim 3.
For all $i\in[n]$, we have $\sum_{j\in[m]}d_{ij}x^{t-1}_{ij}=1/\beta^{t-1}_{i}\cdot(\sum_{j\in[m]}p^{t-1}_{j}x^{t-1}_{ij})$. Analogously, we have $\sum_{j\in[m]}d_{ij}x^{t}_{ij}=1/\beta^{t}_{i}\cdot(\sum_{j\in[m]}p^{t}_{j}x^{t}_{ij})$.
Proof.
Under the complementary slackness condition, we have $x^{t-1}_{ij}(p^{t-1}_{j}-\beta^{t-1}_{i}d_{ij})=0$ for all $i,j$. Therefore, for each $i\in[n]$, we have
$\sum_{j\in[m]}x^{t-1}_{ij}(p^{t-1}_{j}-\beta^{t-1}_{i}d_{ij})=0$
$\implies\sum_{j\in[m]}p^{t-1}_{j}x^{t-1}_{ij}=\beta^{t-1}_{i}\sum_{j\in[m]}d_{ij}x^{t-1}_{ij}$
$\implies\sum_{j\in[m]}d_{ij}x^{t-1}_{ij}=1/\beta^{t-1}_{i}\cdot(\sum_{j\in[m]}p^{t-1}_{j}x^{t-1}_{ij})$
The proof for the second statement of the theorem follows analogously. ∎
From 3, it follows that $\prod_{i\in[n]}(\sum_{j\in[m]}d_{ij}x^{t-1}_{ij})^{B_{i}}=(\prod_{i\in[n]}(1/\beta^{t-1}_{i})^{B_{i}})\cdot(\prod_{i\in[n]}(\sum_{j\in[m]}$ $p^{t-1}_{j}x^{t-1}_{ij})^{B_{i}})$ and 2 it follows that $\prod_{i\in[n]}(\sum_{j\in[m]}d_{ij}x^{t}_{ij})^{B_{i}}=\prod_{i\in[n]}(B_{i}/\beta^{t-1}_{i})^{B_{i}}$ Therefore, if we comparing the weighted product of disutilities in the current iteration and the previous iteration, we have
$\frac{\prod_{i\in[n]}(\sum_{j\in[m]}d_{ij}x^{t}_{ij})^{B_{i}}}{\prod_{i\in[n]}(\sum_{j\in[m]}d_{ij}x^{t-1}_{ij})^{B_{i}}}$
$=$ $\frac{\prod_{i\in[n]}(\frac{B_{i}}{\beta^{t-1}_{i}})^{B_{i}}}{(\prod_{i\in[n]}(\frac{1}{\beta^{t-1}_{i}})^{B_{i}})\cdot(\prod_{i\in[n]}(\sum_{j\in[m]}p^{t-1}_{j}x^{t-1}_{ij})^{B_{i}})}$
$=$ $\frac{\prod_{i\in[n]}B_{i}^{B_{i}}}{\prod_{i\in[n]}(\sum_{j\in[m]}p^{t-1}_{j}x^{t-1}_{ij})^{B_{i}}}$
We now make an observation on $\sum_{i\in[n]}x^{t}_{ij}$– interpreting the dual variables $x^{t}_{ij}$ as the fraction of chore $j$ allocated to agent $i$, one can interpret $\sum_{i\in[n]}x^{t}_{ij}$ as the total consumption of chore $j$ at the end of the $t$-th iteration.
Claim 4.
For all $i\in[n]$, and for all $t\geq 0$, we have $\sum_{i\in[n]}x^{t}_{ij}=\frac{1}{B}\cdot\sum_{i\in[n]}B_{i}\cdot\frac{\beta^{t}_{i}}{\beta^{t-1}_{i}}$.
Proof.
Under stationarity conditions, we have $\sum_{i\in[n]}x^{t}_{ij}=-\delta^{t}$ for all $j$ such that $p^{t}_{j}>0$. Furthermore, from 3, it follows that $\sum_{j\in[m]}p^{t}_{j}x^{t}_{ij}=\beta^{t}_{i}\cdot(\sum_{j\in[m]}d_{ij}x^{t}_{ij})$. Observe that,
$\sum_{i\in[n]}B_{i}\cdot\frac{\beta^{t}_{i}}{\beta^{t-1}_{i}}$ $=\sum_{i\in[n]}\beta^{t}_{i}\cdot\frac{B_{i}}{\beta^{t-1}_{i}}$
$=\sum_{i\in[n]}\beta^{t}_{i}\cdot(\sum_{j\in[m]}d_{ij}x^{t}_{ij})$ (by 2)
$=\sum_{i\in[n]}\sum_{j\in[m]}p^{t}_{j}x^{t}_{ij}$ (by 3)
$=\sum_{j\in[m]}p^{t}_{j}\cdot\sum_{i\in[n]}x^{t}_{ij}$
$=\sum_{j\mid p^{t}_{j}>0}p^{t}_{j}\cdot\sum_{i\in[n]}x^{t}_{ij}$
$=\sum_{j\mid p^{t}_{j}>0}p^{t}_{j}\cdot(-\delta^{t})$
$=(-\delta^{t})\cdot\sum_{j\in[m]}p^{t}_{j}$
$=-(\sum_{i\in[n]}B_{i})\delta^{t}=-B\delta^{t}$
Thus, we have $-\delta^{t}=\frac{1}{B}\cdot\sum_{i\in[n]}B_{i}\cdot\frac{\beta^{t}_{i}}{\beta^{t-1}_{i}}$. Therefore, for all $j\in[m]$, such that $p^{t}_{j}>0$, we have $\sum_{i\in[n]}x^{t}_{ij}=-\delta^{t}=\frac{1}{B}\cdot\sum_{i\in[n]}B_{i}\cdot\frac{\beta^{t}_{i}}{\beta^{t-1}_{i}}$. It only suffices to show that $p^{t}_{j}>0$ for all $j\in[m]$.
We prove this by contradiction. Assume there exists a $j\in[m]$, such that $p^{t}_{j}=0$. Then note that $p^{t}_{j}/d_{ij}<\beta^{t}_{i}$ for all $i\in[n]$ (follows from the fact that every $\beta^{t}_{i}$ is strictly positive). Note that there exists a $\gamma\ll\min_{\ell\in[n]}d_{\ell j}\beta^{t}_{\ell}$, such that if we increase $p^{t}_{j}$ to $\gamma$, and decrease all other prices by a multiplicative factor of $(B-\gamma)$, we still have a feasible set of prices [^9] All the prices are still non-negative and they still sum up to $B$. and $\max_{j\in[m]}p^{t}_{j}/d_{ij}$ strictly decreases for all $i\in[n]$. Let $\beta^{\prime}_{i}=\max_{j\in[m]}p^{t}_{j}/d_{ij}$ (post the multiplicative scaling). Note that $\beta^{\prime}_{i}<\beta^{t}_{i}$ for all $i\in[n]$ and thus $\sum_{i\in[n]}B_{i}\cdot\beta^{\prime}_{i}/\beta^{t-1}_{i}<\sum_{i\in[n]}B_{i}\cdot\beta^{t}_{i}/\beta^{t-1}_{i}$ which contradicts the optimality of $\beta^{t}$. ∎
Claim 5.
If $(p^{t-1},x^{t-1})$ is not an $\epsilon$-approximate equilibrium, then $\log\Bigg{(}\frac{\prod_{i\in[n]}B_{i}^{B_{i}}}{\prod_{i\in[n]}(\sum_{j\in[m]}p^{t-1}_{j}x^{t-1}_{ij})^{B_{i}}}\Bigg{)}\geq\min_{i}B_{i}\cdot\epsilon^{2}/2.3$.
Proof.
Note that $x^{t-1}_{ij}>0$ only if $d_{ij}/p_{j}$ is minimum for all $j\in[m]$ for agent $i$. Thus, if $(p^{t-1},x^{t-1})$ is not an $\epsilon$-approximate equilibrium, then there exists a $j\in[m]$ such that $\sum_{i\in[n]}x^{t-1}_{ij}<1-\epsilon$, or there exists an $i\in[n]$, such that $|\sum_{j\in[m]}p^{t-1}_{j}x^{t-1}_{ij}-B_{i}|>B_{i}\epsilon$.
  • Case $\sum_{i\in[n]}x^{t-1}_{ij}<1-\epsilon$: From 4, it follows that $\sum_{i\in[n]}x^{t-1}_{ij}=\frac{1}{B}\cdot\sum_{i\in[n]}B_{i}\frac{\beta^{t-1}_{i}}{\beta^{t-2}_{i}}<1-\epsilon$, implying that $\sum_{i\in[n]}B_{i}\frac{\beta^{t-1}_{i}}{\beta^{t-2}_{i}}<B(1-\epsilon)$. Further, observe that
    $\sum_{j\in[n]}p^{t-1}_{j}\cdot x^{t-1}_{ij}$ $=\beta^{t-1}_{i}\cdot\sum_{j\in[m]}d_{ij}x^{t-1}_{ij}$ (by 3)
    $=\beta^{t-1}_{i}\cdot\frac{B_{i}}{\beta^{t-2}_{i}}$ (by 2)
    Thus, we have $\sum_{i\in[n]}\sum_{j\in[m]}p^{t-1}_{j}x^{t-1}_{ij}=\sum_{i\in[n]}B_{i}\frac{\beta^{t-1}_{i}}{\beta^{t-2}_{i}}<B(1-\epsilon)$. This implies that $\prod_{i\in[n]}(\sum_{j\in[m]}$ $p^{t-1}_{j}x^{t-1}_{ij})^{B_{i}}\leq\prod_{i\in[n]}(B_{i}(1-\epsilon))^{B_{i}}$, further implying that
    $\frac{\prod_{i\in[n]}B_{i}^{B_{i}}}{\prod_{i\in[n]}(\sum_{j\in[m]}p^{t-1}_{j}x^{t-1}_{ij})^{B_{i}}}$ $\geq\frac{1}{(1-\epsilon)^{B}}$
    $\geq(1+\epsilon)^{B}$
    $\geq 1+{B\epsilon}$
    Therefore, $\log\Bigg{(}\frac{\prod_{i\in[n]}B_{i}^{B_{i}}}{\prod_{i\in[n]}(\sum_{j\in[m]}p^{t-1}_{j}x^{t-1}_{ij})^{B_{i}}}\Bigg{)}\geq\log(1+B\epsilon)\geq\min_{i}B_{i}\epsilon^{2}/2.3$ as $\log(1+x)\geq x^{2}/2.3$ for all $x\in[0,1]$.
  • Case $|\sum_{j\in[m]}p^{t-1}_{j}x^{t-1}_{ij}-B_{i}|>B_{i}\epsilon$: Let $i$ be an agent such that $\sum_{j\in[m]}p^{t-1}_{j}x^{t-1}_{ij}=B_{i}(1+\epsilon^{\prime})$, where $\epsilon^{\prime}\in\{-\theta,\theta\}$ for some $\theta>\epsilon$. Observe that $\sum_{\ell\in[n]\setminus\{i\}}(\sum_{j\in[m]}p^{t-1}_{j}x^{t-1}_{\ell j})<B-B_{i}-B_{i}\epsilon^{\prime}$, as $\sum_{i\in[n]}(\sum_{j\in[m]}p^{t-1}_{j}x^{t-1}_{ij})=\sum_{i\in[n]}B_{i}\frac{\beta^{t-1}_{i}}{\beta^{t-2}_{i}}\leq B$, and $\sum_{j\in[m]}p^{t-1}_{j}x^{t-1}_{ij}=B_{i}(1+\epsilon^{\prime})$. This implies that $\prod_{\ell\in[n]\setminus\{i\}}(\sum_{j\in[m]}p^{t-1}_{j}x^{t-1}_{\ell j})^{B_{i}}\leq\prod_{\ell\in[n]\setminus\{i\}}(B_{\ell}\cdot(1-\frac{B_{i}\epsilon^{\prime}}{B-B_{i}}))^{B_{\ell}}$. Before, we proceed, we mention the following useful fact,
    Fact 1.
    For all real $x$, we have, $\log(e^{x}/(1+x))\geq x^{2}/2.3$ for all $x\in[-1,1]$.
    Proof.
    We have,
    $\hskip 20.0pt\frac{\prod_{i\in[n]}B_{i}^{B_{i}}}{\prod_{i\in[n]}(\sum_{j\in[m]}p^{t-1}_{j}x^{t-1}_{ij})^{B_{i}}}$
    $>\frac{\prod_{i\in[n]}B_{i}^{B_{i}}}{(B_{i}(1+\epsilon^{\prime}))^{B_{i}}\prod_{\ell\in[n]\setminus\{i\}}(B_{\ell}\cdot(1-\frac{B_{i}\epsilon^{\prime}}{B-B_{i}}))^{B_{\ell}}}$
    $=\frac{1}{(1+\epsilon^{\prime})^{B_{i}}(1-\frac{B_{i}\epsilon^{\prime}}{B-B_{i}})^{(B-B_{i})}}$
    $\geq\frac{1}{(1+\epsilon^{\prime})^{B_{i}}e^{-\frac{B_{i}\epsilon^{\prime}}{B-B_{i}}\cdot(B-B_{i})}}$
    $=\frac{1}{(1+\epsilon^{\prime})^{B_{i}}e^{-B_{i}\epsilon^{\prime}}}$
    $=\bigg{(}\frac{1}{(1+\epsilon^{\prime})e^{-\epsilon^{\prime}}}\bigg{)}^{B_{i}}$
Now, observe that
$\log\Bigg{(}\frac{\prod_{i\in[n]}B_{i}^{B_{i}}}{\prod_{i\in[n]}(\sum_{j\in[m]}p^{t-1}_{j}x^{t-1}_{ij})^{B_{i}}}\Bigg{)}$ $>\log\Bigg{(}\bigg{(}\frac{1}{(1+\epsilon^{\prime})e^{-\epsilon^{\prime}}}\bigg{)}^{B_{i}}\Bigg{)}$
$=B_{i}\log\Bigg{(}\frac{e^{\epsilon^{\prime}}}{1+\epsilon^{\prime}}\Bigg{)}$
$\geq B_{i}\frac{\epsilon^{\prime 2}}{2.3}$ (from Fact 1)
$=\min_{i}B_{i}\frac{\theta^{2}}{2.3}=\min_{i}B_{i}\frac{\epsilon^{2}}{2.3}.$
We are now ready to prove our main result.
Theorem 8.
Let $T\geq\frac{2.3(f(\beta^{*})-f(\beta^{0}))}{\epsilon^{2}\min_{i}B_{i}}$. Then there exists a $t\leq T$, such that $(p^{t},x^{t})$ is a $\epsilon$-CE
Proof.
Assume otherwise, i.e., for all $t\in[T]$, $(p^{t},x^{t})$ is not a $\epsilon$-CE. Recall that we have,
$\frac{\prod_{i\in[n]}(\sum_{j\in[m]}d_{ij}x^{t}_{ij})^{B_{i}}}{\prod_{i\in[n]}(\sum_{j\in[m]}d_{ij}x^{t-1}_{ij})^{B_{i}}}=\frac{\prod_{i\in[n]}B_{i}^{B_{i}}}{\prod_{i\in[n]}(\sum_{j\in[m]}p^{t-1}_{j}x^{t-1}_{ij})^{B_{i}}}$
By 2, we can replace $\sum_{j\in[m]}d_{ij}x^{t}_{ij}$ with $1/\beta^{t-1}_{i}$, implying,
$\frac{\prod_{i\in[n]}(1/\beta^{t-1}_{i})^{B_{i}}}{\prod_{i\in[n]}(1/\beta^{t-2}_{i})^{B_{i}}}=\frac{\prod_{i\in[n]}B_{i}^{B_{i}}}{\prod_{i\in[n]}(\sum_{j\in[m]}p^{t-1}_{j}x^{t-1}_{ij})^{B_{i}}}$ (20)
Since $(p^{t-1},x^{t-1})$ is not a $\epsilon$-CEEI, we have $\log\bigg{(}\frac{\prod_{i\in[n]}B_{i}^{B_{i}}}{\prod_{i\in[n]}(\sum_{j\in[m]}p^{t-1}_{j}x^{t-1}_{ij})^{B_{i}}}\bigg{)}\geq\min_{i}B_{i}\cdot\epsilon^{2}/2.3$ by 5. Taking logs on both sides of Eq. 20, we have
$f(\beta^{t-1})-f(\beta^{t-2})\geq\min_{i}B_{i}\cdot\epsilon^{2}/2.3$ (21)
Summing up Eq. 21 from $t=2$ to $T+1$, we have $f(\beta^{T})-f(\beta^{0})>\min_{i}B_{i}\cdot T\epsilon^{2}/2.3$. Since $f(\beta^{*})\geq f(\beta^{T})$, this implies that $f(\beta^{*})-f(\beta^{0})>\min_{i}B_{i}\cdot T\epsilon^{2}/2.3$, implying that $T<2.3(f(\beta^{*})-f(\beta^{0}))/(\epsilon^{2}\cdot\min_{i}B_{i})$, which is a contradiction. ∎
Remark.
This analysis also gives us additional insights: for instance, any iteration where the solution of the GFW is not a good approximation of a CE, will see a large improvement in the objective function, implying that GFW iterations with bad approximate solutions are small.
Appendix B Approximate CE and Exact CE can be far Apart
We show that approximate CE and Exact CE can be far apart. Consider the following chores market: 2 agents $a_{1}$ and $a_{2}$, and 2 chores $c_{1}$ and $c_{2}$. Set $d_{11}=1$, and $d_{12}=M$ for $M\gg 1$. Set $d_{21}=1-\epsilon$, and $d_{22}=1+\epsilon$ for $\epsilon\ll 1$. It is easy to verify that the only CEEI here is
  • $p_{1}=2/(M+1)$ and $p_{2}=2M/(M+1)$,
  • $x_{11}=1$, $x_{21}=(M-1)/(2M)$, and
  • $x_{21}=0$, and $x_{22}=(M+1)/(2M)$.
Note that $\beta_{1}=2/(M+1)$, and $\beta_{2}=2M/((M+1)(1+\epsilon))$. However, the following is an $\epsilon$-approximate CEEI, with $\beta_{1}=\beta_{2}=1$, which is far from the only exact CEEI!
  • $p_{1}=1-\epsilon$ and $p_{2}=1+\epsilon$,
  • $x_{11}=1-\epsilon$, $x_{21}=0$, and
  • $x_{21}=0$, and $x_{22}=1$.
Appendix C Additional numerical experiments
uniform(0,1) Refer to caption Refer to caption Refer to caption lognormal Refer to caption Refer to caption Refer to caption truncated normal Refer to caption Refer to caption Refer to caption exponential Refer to caption Refer to caption Refer to caption randint(1,1000) Refer to caption Refer to caption Refer to caption
Figure 5: Numerical results for 100 buyers and up to 1000 chores. GFW solves every instance for every distribution.
uniform(0,1) Refer to caption Refer to caption Refer to caption lognormal Refer to caption Refer to caption Refer to caption truncated normal Refer to caption Refer to caption Refer to caption exponential Refer to caption Refer to caption Refer to caption randint(1,1000) Refer to caption Refer to caption Refer to caption
Figure 6: Numerical results with very small Gurobi tolerance parameters.